The measured strain values at point Q are as follows: Ea = 40(10), Eb = 980(10), &c = 330(10) 1) Calculate the strain compone《 s,
Answer #1

As per the question: . - 40х10; S400x106 , - 330 x 106 oe koo : a = + + - со 20 + лау. So20 -0 since a is parallel to t-axisS 20 t Sen 20 - bis at angle of 0= 60 7 Ep. Em 8 + Sqz Ey cas 20 + TV, 5.020 980 %10 €Ent + Basen Cos 128 + mogu so izá 960 x. ЗАКОХТО 40 хв“х4 2 + 2 Ey - E, + y + (3 ху » E +38 + 2 = 3420 x 10 € 3 + 3 y = 3920 –40) нь с 3€+ з ту . 38 от ог“ – 1. c iAdding a ☺ & ③ fEy = 5160x106 - Ey = 560X1697 from a 3 Vay: 3x8 – 1280 x10° Ney . (2550 - 1280) x106 - V3 Nay = 750.55 x106)E nay Trys G. Ney = 2(17) 4 : 16.4 x 103x 70. KT x 06 Jay = 57.72 mail 3. Principal strenes: Ce sit nya tahun i 2 65.16 4190.